80 people showed up for a party, 50 were men, what % were women?

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

---
(80-50)/80 * 100 = 37.5%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

80 people showed up for a party, 50 were men, what % were women?
-----
# of women 80-50 = 30
-----
Percent of women = 30/80 = 3/8 = 37.5%
-----------------------------------------
Cheers,
Stan H.
----------------